Operating Systems and Systems Integration

Burning Red Hat CDs with our CD Writer 1 Aim

We have recently installed a RICOH CD writer on a PII in A362. So far, only the Linux software has been installed. This is a quick guide to using it to burn Linux CDs.

2 Background

Nick Urbanik <nicku@nicku.org>  ver. 1.1